# Graphon Quantum Filtering systems

### v1

@@ 5.2 Propagation of chaos
$\mathbb{E}\Big{[}\Big{\|}A_{ji_{l}}\hat{\tau_{1}h_{i}}\hat{\tau_{1}h_{j}}\boldsymbol{\gamma}_{t,j}\boldsymbol{\gamma}_{t,i_{l}}\Psi_{t}^{cN}\Big{\|}_{2}^{2}\Big{]}$ $\leq\mathbb{E}\Big{[}\langle\Psi_{t}^{cN},\hat{\tau_{1}h_{i}}\hat{\tau_{1}h_{j}}\boldsymbol{\gamma}_{t,i_{l}}\boldsymbol{\gamma}_{t,j}A_{ji_{l}}^{2}\hat{\tau_{1}h_{i}}\hat{\tau_{1}h_{j}}\boldsymbol{\gamma}_{t,i_{l}}\boldsymbol{\gamma}_{t,j}\Psi_{t}^{cN}\rangle\Big{]}$
$\leq\|a\|_{2}\mathbb{E}\Big{[}\langle\Psi_{t}^{cN},\hat{\tau_{1}h_{i}^{2}}\Psi_{t}^{cN}\rangle\langle\Psi_{t}^{cN},\hat{\tau_{1}h_{j}^{2}}\Psi_{t}^{cN}\rangle\Big{]}$
$\leq\|a\|_{2}\mathbb{E}\Big{[}\Big{(}\langle\Psi_{t}^{cN},\hat{f_{i}}\Psi_{t}^{cN}\rangle+\frac{1}{N}\Big{)}\Big{(}\langle\Psi_{t}^{cN},\hat{f_{j}}\Psi_{t}^{cN}\rangle+\frac{1}{N}\Big{)}\Big{]}$
$\leq\|a\|_{2}\mathbb{E}\Big{[}\Big{(}\mathcal{D}^{i}_{N}(t)+\frac{1}{N}\Big{)}\Big{(}\mathcal{D}^{j}_{N}(t)+\frac{1}{N}\Big{)}\Big{]}$
$\leq\|a\|_{2}\mathbb{E}\Big{[}\Big{(}\mathcal{D}_{N}(t)+\frac{1}{N}\Big{)}^{2}\Big{]}\leq\|a\|_{2}\underbrace{\mathbb{E}\Big{[}\Big{(}\mathcal{D}_{N}(t)+\frac{3}{N}\Big{)}\Big{]}}_{{y}}.$
Hence,
$\mathbb{E}\Big{[}(\triangle_{2}^{j,i_{l}})\Big{]}$ $\stackrel{{\scriptstyle\sqrt{xy}\leq\frac{x+y}{2}}}{{\leq}}C\Big{(}\mathbb{E}\Big{[}\mathcal{D}_{N}(t)\Big{]}+\frac{3}{2N}\Big{)}.$
Estimate $(\triangle_{3}^{j,i_{l}})$
We have
$\mathbb{E}\big{[}{(\triangle_{3}^{j,i_{l}})}\big{]}$ $=\mathbb{E}\Big{[}\Big{|}\langle\Psi_{t}^{cN},\boldsymbol{\gamma}_{t,j}\boldsymbol{g}_{t,i_{l}}\big{(}A_{ji_{l}}-A^{\gamma_{t,i_{l}}}_{j}\big{)}\boldsymbol{g}_{t,j}\boldsymbol{g}_{t,i_{l}}\Psi_{t}^{cN}\rangle\Big{|}\Big{]}$
$\stackrel{{\scriptstyle\text{C-S}}}{{\leq}}\mathbb{E}\Big{[}\Big{\|}\boldsymbol{g}_{t,j}\boldsymbol{g}_{t,i_{l}}\Psi_{t}^{cN}\Big{\|}_{2}\Big{\|}\big{(}A_{ji_{l}}-A^{\gamma_{t,i_{l}}}_{j}\big{)}\boldsymbol{\gamma}_{t,j}\boldsymbol{g}_{t,i_{l}}\Psi_{t}^{cN}\Big{\|}_{2}\Big{]}$
$\leq\mathbb{E}\Big{[}\Big{\|}\boldsymbol{g}_{t,j}\boldsymbol{g}_{t,i_{l}}\Psi_{t}^{cN}\Big{\|}_{2}\Big{\|}\boldsymbol{\gamma}_{t,j}\big{(}A_{ji_{l}}-A^{\gamma_{t,i_{l}}}_{j}\big{)}^{2}\boldsymbol{\gamma}_{t,j}\Psi_{t}^{cN}\Big{\|}_{\text{op}}\Big{\|}\boldsymbol{g}_{t,i_{l}}\Psi_{t}^{cN}\Big{\|}_{2}\Big{]}$
$\leq C\mathbb{E}\Big{[}\mathcal{D}_{N}^{i_{l}}(t)+\mathcal{D}_{N}^{j_{1}}{t})\Big{]}\leq C\mathbb{E}\Big{[}\mathcal{D}_{N}(t)\Big{]},$
where we used the bound $\Big{\|}\boldsymbol{\gamma}_{t,j}\big{(}A_{ji_{l}}-A^{\gamma_{t,i_{l}}}_{j}\big{)}^{2}\boldsymbol{\gamma}_{t,j}\Psi_{t}^{cN}\Big{\|}_{\text{op}}\leq C.$ We thus have
$\mathbb{E}\Big{[}\Big{|}(\star\star\star)_{12}^{j_{1}}\Big{|}\Big{]}$ $\leq C\big{(}\mathbb{E}[\mathcal{D}_{N}(t)]+\frac{1}{N}\big{)},$
and,
$\mathbb{E}\Big{[}\Big{|}(\star\star\star)_{1}^{j_{1}}\Big{|}\Big{]}$ $\leq\mathbb{E}\Big{[}\Big{|}(\star\star\star)_{11}^{j_{1}}\Big{|}+\mathbb{E}\Big{[}\Big{|}(\star\star\star)_{12}^{j_{1}}\Big{|}\leq C\big{(}\mathbb{E}[\mathcal{D}_{N}(t)]+\frac{1}{\sqrt{N}}\big{)}.$
Estimate $(\star\star\star)_{2}^{j_{1}}$
@@
$\mathbb{E}\Big{[}\Big{|}(\star\star\star)_{2}^{j_{1}}\Big{|}\Big{]}$ $\leq C\big{(}\mathbb{E}[\mathcal{D}_{N}(t)]+\frac{1}{\sqrt{N}}\big{)}.$
@@
$\mathrm{d}\mathbb{E}[\mathcal{D}_{N}^{j}(t)]$ $\leq C\Big{(}\|w^{cN}-w\|_{\square}+\frac{1}{\sqrt{N}}+\mathbb{E}\big{[}\mathcal{D}_{N}(t)\big{]}\Big{)}.$
@@
$\mathrm{d}\mathbb{E}[\mathcal{D}_{N}(t)]$ $\leq C\Big{(}\|w^{cN}-w\|_{\square}+\frac{1}{\sqrt{N}}+\mathbb{E}\big{[}\mathcal{D}_{N}(t)\big{]}\Big{)},$
@@ 6 Application
In this section, we examine how the graphon quantum filtering system can be used for application. To avoid unnecessary technical difficulties, we restrict to a finite-dimensional Hilbert space $\mathbb{H}=\mathbb{C}^{d}$.
@@ 6.1 Graphon quantum dynamic game
$\quad(s,\rho)\longrightarrow{\alpha}(s,\rho)$
@@
Given an admissible control strategy $\underline{\alpha}$, the state of the controlled graphon quantum filtering system evolves as:
$\mathrm{d}\gamma_{t,u}^{\underline{\alpha}}$ $=-\mathrm{i}\Big{[}\tilde{H}+\alpha_{u}(t)\hat{H}+\int_{0}^{1}w(u,v)A^{\mathbb{E}[\gamma_{t,v}^{\underline{\alpha}}]}\mathrm{d}v,\gamma_{t,u}^{\underline{\alpha}}\Big{]}\mathrm{d}t+\Big{(}L\gamma_{t,u}^{\underline{\alpha}}L^{\dagger}-\frac{1}{2}\{L^{\dagger}L,\gamma_{t,u}^{\underline{\alpha}}\}\Big{)}\mathrm{d}t$
@@
$\mathcal{J}_{u}(\beta,\Gamma_{u}^{\underline{\alpha}}):=\mathbb{E}\Big{[}\int_{0}^{T}C\big{(}\beta,\gamma_{s,u}^{(\beta,\underline{\alpha}_{-u})},\Gamma_{s,u}^{\underline{\alpha}}\big{)}\mathrm{d}s+F(\gamma_{T}^{\underline{\alpha}},\Gamma_{T,u}^{\underline{\alpha}})\Big{]},$
@@
The measurable functions $C:\mathbb{R}\times\mathcal{M}_{d}(\mathbb{C})\times\mathcal{M}_{d}(\mathbb{C})\to\mathbb{R},\;\text{and \;}F:\mathcal{M}_{d}(\mathbb{C})\times\mathcal{M}_{d}(\mathbb{C})\to\mathbb{R}$ are the current and terminal cost, respectively. The graphon-weighted aggregate $\Gamma_{u}^{\underline{\alpha}}:=(\Gamma_{t,u}^{\underline{\alpha}})_{0\leq t\leq T}$ is defined by
$\Gamma_{t,u}^{\underline{\alpha}}:=\int_{0}^{1}w(u,v)\gamma_{t,v}^{\underline{\alpha}}\lambda(\mathrm{d}v),\quad\forall t\in[0,T].$
We now adopt the following important notion of Nash equilibrium,
Definition 4.
@@
$\mathcal{J}_{u}(\hat{\alpha}_{u},\Gamma_{u}^{\underline{\hat{\alpha}}})$ $\leq\mathcal{J}_{u}(\beta,\Gamma_{u}^{\underline{\hat{\alpha}}}),\quad\beta\in\mathcal{U}(u),\;\forall u\in I.$
@@
The question of the existence of equilibria can be addressed using classical arguments from stochastic optimal control, such as the stochastic maximum principle and dynamic programming principle; see e.g., [71, 17]. We leave this and related issues to future work.
6.2 Quantum state preparation
An important objective in quantum control theory is quantum state preparation which is addressed in many papers, see e.g., [65, 51, 48, 22, 16].
In this section, we consider a concrete example of a graphon-qubit system. Fix the parameters as follows:
$\eta$ $=1,\;\;w(u,v)=uv,\;\;\gamma_{0,u}=\begin{pmatrix}\frac{1}{2}&0\\ 0&\frac{1}{2}\end{pmatrix},$
$\mathcal{J}_{-}$ $=[0,\frac{1}{2}),\;\;\mathcal{J}_{+}=[\frac{1}{2},1].$
Let’s set Pauli matrices and ground and excited states of a two-level system,
$\rho_{g}:=\begin{pmatrix}0&0\\ 0&1\end{pmatrix},\;\rho_{e}:=\begin{pmatrix}1&0\\ 0&0\end{pmatrix},\;\sigma_{x}:=\begin{pmatrix}0&-1\\ 1&0\end{pmatrix},\;\sigma_{z}:=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$
Schematic picture : Graphon $w(u,v)=uv.$
Under an admissible control field $\underline{\alpha}=\bigl{(}\alpha_{u}(t)\bigr{)}_{u\in[0,1]}$, the graphon Belavkin system evolves, for all $u\in[0,1]$ and a.s. for $t\geq 0$, according to:
$\mathrm{d}\gamma_{t,u}$ $=-\mathrm{i}\bigl{[}\alpha_{u}(t)\sigma_{x}+u\Bigl{(}\int_{0}^{1}v\mathbb{E}[z_{t,v}]\mathrm{d}v\Bigr{)}\sigma_{z},\gamma_{t,u}\bigr{]}\mathrm{d}t+\bigl{(}\sigma_{z}\gamma_{t,u}\sigma_{z}-\gamma_{t,u}\bigr{)}\mathrm{d}t$ (6.1)
$z_{t,u}:=\mathrm{t}\mathrm{r}(\gamma_{t,u}\sigma_{z}).$
Quantum state reduction
Setting $\underline{\alpha}\equiv 0$ turns into an uncontrolled dynamics. Define the stochastic Lyapunov function as $(V_{t})_{t\geq 0}$
$V_{t,u}:=1-z_{t,u}^{2},\ V_{t}:=\int_{I}V_{t,u}\mathrm{d}u.$
Applying Itô’s formula, we obtain
$\mathrm{d}z_{t,u}^{2}$ $=2z_{t,u}\mathrm{d}z_{t,u}+4(1-z_{t,u}^{2})^{2}\mathrm{d}t$
$=4(1-z_{t,u}^{2})^{2}\mathrm{d}t+2z_{t,u}(1-z_{t,u}^{2})\mathrm{d}W_{t}^{u},$
so that
Taking the expectation, we have
$\mathbb{E}[V_{t}]$ $=V_{0}-4\int_{0}^{t}\mathbb{E}[V_{s}^{2}]\mathrm{d}s.$
Since $0\leq V_{t}\leq 1$, it follows that a.s. $\int_{0}^{\infty}V_{s}^{2}\mathrm{d}s<\infty$. As $V_{t}$ is a non-negative supermartingale, bounded below and with decreasing expectation, we conclude that a.s. $V_{t}\to 0$. Thus, $V_{t,u}\to 0$ for all $u\in I$, which implies that $\gamma_{t,u}\to\{\rho_{e},\rho_{g}\}$. Hence, each particles states converges asymptotically to one of the projectors $\rho_{g}$ or $\rho_{e}$.
Deterministic state preparation
For $u\in\mathcal{J}\pm$, we adapt the control strategy from [69] and set:
$\alpha_{u}(t)=\begin{cases}-8\mathrm{i}\mathrm{t}\mathrm{r}\bigl{(}[\sigma_{x},\gamma_{t,u}]\rho_{g}\bigr{)}+5\bigl{(}1-\mathrm{t}\mathrm{r}(\gamma_{t,u}\rho_{g})\bigr{)},&u\in\mathcal{J}_{-},\\ -8\mathrm{i}\mathrm{t}\mathrm{r}\bigl{(}[\sigma_{x},\gamma_{t,u}]\rho_{e}\bigr{)}+5\bigl{(}1-\mathrm{t}\mathrm{r}(\gamma_{t,u}\rho_{e})\bigr{)},&u\in\mathcal{J}_{+}.\end{cases}$ (6.2)
Under this feedback control, state preparation becomes deterministic in the sense that particles with $u\in\mathcal{J}_{-}$ converge to the ground state $\rho_{g}$, while particles with $u\in\mathcal{J}_{+}$ converge to the excited state $\rho_{e}$.
Simulations
We simulate the system (6.1) in two scenarios: the uncontrolled case with $\underline{\alpha}\equiv 0$ and the controlled case using the strategy defined in (6.2). In both cases, we plot the fidelity $\mathcal{F}$, defined by $\mathcal{F}(\gamma,\rho):=\big{(}\mathrm{t}\mathrm{r}(\sqrt{\sqrt{\rho}\gamma\sqrt{\rho}})\big{)}^{2},$ between the quantum state $\gamma_{t,u}$ and the ground state $\rho_{g}$ for particles $u\in\mathcal{J}_{-}$ (in red) and $u\in\mathcal{J}_{+}$ (in blue).
The fidelity satisfies:
Figure 1 illustrates the case of uncontrolled quantum state reduction, while Figure 2 shows the outcome under the control law (6.2).
Figure 1: Quantum state reduction under $\underline{\alpha}\equiv 0$.
Figure 2: Quantum state preparation using the control law (6.2).

### v3

@@ 5.2 Propagation of chaos
$\mathbb{E}_{\mathbb{P}}\Big[\Big|(\star\star\star)_{1}^{j_{1}}\Big|\Big]$ $\leq\mathbb{E}_{\mathbb{P}}\Big[\Big|(\star\star\star)_{11}^{j_{1}}\Big|\Big]+\mathbb{E}_{\mathbb{P}}\Big[\Big|(\star\star\star)_{12}^{j_{1}}\Big|\Big]\leq C\Big(\mathbb{E}_{\mathbb{P}}[\mathcal{D}_{N}(t)]+\frac{1}{\sqrt{N}}\Big).$
Estimate $(\star\star\star)_{2}^{j_{1}}$.
@@
$\mathbb{E}_{\mathbb{P}}\Big[\Big|(\star\star\star)_{2}^{j_{1}}\Big|\Big]$ $\leq C\Big(\mathbb{E}_{\mathbb{P}}[\mathcal{D}_{N}(t)]+\frac{1}{\sqrt{N}}\Big).$
@@
$\mathrm{d}\mathbb{E}_{\mathbb{P}}[\mathcal{D}_{N}^{j}(t)]$ $\leq C\Big(\|w^{cN}-w\|_{\square}+\frac{1}{\sqrt{N}}+\mathbb{E}_{\mathbb{P}}\big[\mathcal{D}_{N}(t)\big]\Big).$
@@
$\mathrm{d}\mathbb{E}_{\mathbb{P}}[\mathcal{D}_{N}(t)]$ $\leq C\Big(\|w^{cN}-w\|_{\square}+\frac{1}{\sqrt{N}}+\mathbb{E}_{\mathbb{P}}\big[\mathcal{D}_{N}(t)\big]\Big),$
@@ 6 Application
In this section, we explore potential applications of the graphon quantum filtering system. To keep the presentation simple and avoid unnecessary technicalities, we restrict attention to a finite-dimensional Hilbert space $\mathbb{H}=\mathbb{C}^{d}$. In this case, $\mathcal{B}(\mathbb{H})=\mathcal{M}_{d}(\mathbb{C})$, the set of complex $d\times d$ matrices, which is isomorphic to $\mathbb{R}^{2d^{2}}$. Hence, the spaces can be identified as the standard Euclidean spaces.
6.1 Quantum state preparation
An important objective in quantum control theory is quantum state preparation, which has been studied extensively, see e.g., [71, 56, 68, 18, 52, 24, 4]. This task is challenging due to the high dimensionality of the system, and various methods have been developed to manage this complexity (see e.g., [78, 75, 39]). Here we use the formalism of the mean-field limiting graphon as an approximation to the $N$-body system.
As an illustration, we consider a concrete example of a graphon–qubit system with Ising-type interactions.
We begin by recalling the Pauli matrices and the ground and excited states of a two-level system:
$\rho_{g}:=\begin{pmatrix}0&0\\ 0&1\end{pmatrix},\;\rho_{e}:=\begin{pmatrix}1&0\\ 0&0\end{pmatrix},\;\sigma_{x}:=\begin{pmatrix}0&-1\\ 1&0\end{pmatrix},\;\sigma_{z}:=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$.
We fix the parameters of a system as follows:
$\eta$ $=1,\;\;w(u,v)=uv,\;\;\gamma_{0,u}=\begin{pmatrix}\tfrac{1}{2}&0\\ 0&\tfrac{1}{2}\end{pmatrix}$,
$\mathcal{J}_{-}$ $=[0,\frac{1}{2}),\;\;\mathcal{J}_{+}=[\frac{1}{2},1].$
Schematic picture: Graphon $w(u,v)=uv.$
Under an admissible control $\underline{\alpha}(t)=\{\alpha_{u}(t),u\in I\}$, the graphon Belavkin system evolves, for all $u\in I$ and a.s. for $t\geq 0$, according to
$\mathrm{d}\gamma_{t,u}$ $=-\mathrm{i}\bigl[\alpha_{u}(t)\sigma_{x}+u\Bigl(\int_{0}^{1}v\mathbb{E}_{\mathbb{P}}[z_{t,v}]\mathrm{d}v\Bigr)\sigma_{z},\gamma_{t,u}\bigr]\mathrm{d}t+\bigl(\sigma_{z}\gamma_{t,u}\sigma_{z}-\gamma_{t,u}\bigr)\mathrm{d}t$ (15)
$z_{t,u}:=\mathrm{t}\mathrm{r}(\gamma_{t,u}\sigma_{z}).$
6.1.1 Asymptotic behavior without control
Setting $\underline{\alpha}\equiv 0$ turns into an uncontrolled dynamics. We define the stochastic Lyapunov function $(V_{t,u})_{t\geq 0}$ by
$V_{t,u}:=1-z_{t,u}^{2},\;\;\mathcal{V}_{t}(u,\omega):=V_{t,u}(\omega).$
Applying Itô’s formula, we obtain
$\mathrm{d}z_{t,u}^{2}$ $=2z_{t,u}\mathrm{d}z_{t,u}+4(1-z_{t,u}^{2})^{2}\mathrm{d}t$
$=4(1-z_{t,u}^{2})^{2}\mathrm{d}t+2z_{t,u}(1-z_{t,u}^{2})\mathrm{d}W_{t}^{u},$
so that
Taking the expectation, we have
$\mathbb{E}_{\mathbb{P}}[V_{t,u}]$ $=V_{0}-4\int_{0}^{t}\mathbb{E}_{\mathbb{P}}[V_{s,u}^{2}]\mathrm{d}s.$
Then,
$\mathbb{E}_{\mathcal{P}}[\mathcal{V}_{t}]$ $\leq V_{0}-4\int_{0}^{t}\mathbb{E}_{\mathcal{P}}[\mathcal{V}_{s}^{2}]\mathrm{d}s.$
Since $0\leq\mathcal{V}_{t}\leq 1$, it follows that a.s. $\int_{0}^{\infty}\mathcal{V}_{s}^{2}\mathrm{d}s<\infty$. As $\mathcal{V}_{t}$ is a non-negative supermartingale, bounded below and with decreasing expectation, we conclude that a.s. $\mathcal{V}_{t}\to 0$. Thus $V_{t,u}\to 0$ for all $u\in I$, which implies that $\gamma_{t,u}\to\{\rho_{e},\rho_{g}\}$. Hence, the state of each particle asymptotically collapses to one of the projectors $\rho_{g}$ or $\rho_{e}$.
6.1.2 Deterministic state preparation
For $u\in\mathcal{J}_{\pm}$, we adapt the feedback control strategy from [77] and set
$\alpha_{u}(t)=\begin{cases}-8\mathrm{i}\mathrm{t}\mathrm{r}\bigl([\sigma_{x},\gamma_{t,u}]\rho_{g}\bigr)+5\bigl(1-\mathrm{t}\mathrm{r}(\gamma_{t,u}\rho_{g})\bigr),&u\in\mathcal{J}_{-},\\ -8\mathrm{i}\mathrm{t}\mathrm{r}\bigl([\sigma_{x},\gamma_{t,u}]\rho_{e}\bigr)+5\bigl(1-\mathrm{t}\mathrm{r}(\gamma_{t,u}\rho_{e})\bigr),&u\in\mathcal{J}_{+}.\end{cases}$ (16)
Under this feedback control, quantum state preparation becomes deterministic in the sense that particles with $u\in\mathcal{J}_{-}$ converge to the ground state $\rho_{g}$, while particles with $u\in\mathcal{J}_{+}$ converge to the excited state $\rho_{e}$.
6.1.3 Simulations
We simulate the dynamics of system (15) under two scenarios: the uncontrolled case with $\underline{\alpha}\equiv 0$ and the controlled case using the feedback strategy defined in (16). In both cases, we evaluate the fidelity $\mathcal{F}$ between the evolving quantum state $\gamma_{t,u}$ and the ground state $\rho_{g}$, defined by
$\mathcal{F}(\gamma,\rho):=\big(\mathrm{t}\mathrm{r}(\sqrt{\sqrt{\rho}\gamma\sqrt{\rho}})\big)^{2}.$
For particles with $u\in\mathcal{J}_{-}$, the fidelity with respect to $\rho_{g}$ is plotted in red, and for those with $u\in\mathcal{J}_{+}$ in blue.
The fidelity satisfies:
Figure 1 illustrates the case of uncontrolled quantum state reduction, while Figure 2 shows the deterministic stabilization achieved under the control law (16).
Figure 1: Asymptotic behavior under $\underline{\alpha}\equiv 0$.
Figure 2: Quantum state preparation using the control law (16).
@@ 6.2 Graphon quantum dynamic game
$\quad(s,\rho)\longrightarrow{\alpha}(s,\rho).$
@@
Given an admissible control strategy $\underline{\alpha}$, the state of the controlled graphon quantum filtering system evolves as
$\mathrm{d}\gamma_{t,u}^{\underline{\alpha}}$ $=-\mathrm{i}\Big[\tilde{H}+\alpha_{u}(t)\hat{H}+\int_{0}^{1}w(u,v)A^{\mathbb{E}_{\mathbb{P}}[\gamma_{t,v}^{\underline{\alpha}}]}\mathrm{d}v,\gamma_{t,u}^{\underline{\alpha}}\Big]\mathrm{d}t+\Big(L\gamma_{t,u}^{\underline{\alpha}}L^{\dagger}-\frac{1}{2}\{L^{\dagger}L,\gamma_{t,u}^{\underline{\alpha}}\}\Big)\mathrm{d}t$
@@
$\mathcal{J}_{u}(\beta,\mathcal{A}_{u}^{\underline{\alpha}}):=\mathbb{E}_{\mathbb{P}}\Big[\int_{0}^{T}C\big(\beta,\gamma_{s,u}^{(\beta,\underline{\alpha}_{-u})},\mathcal{A}_{s,u}^{\underline{\alpha}}\big)\mathrm{d}s+F(\gamma_{T}^{\underline{\alpha}},\mathcal{A}_{T,u}^{\underline{\alpha}})\Big],$
@@
The measurable functions $C:\mathbb{R}\times\mathcal{M}_{d}(\mathbb{C})\times\mathcal{M}_{d}(\mathbb{C})\to\mathbb{R},\;\text{and \;}F:\mathcal{M}_{d}(\mathbb{C})\times\mathcal{M}_{d}(\mathbb{C})\to\mathbb{R}$ are the current and terminal cost, respectively. The graphon-weighted aggregate $\mathcal{A}_{u}^{\underline{\alpha}}:=(\mathcal{A}_{t,u}^{\underline{\alpha}})_{0\leq t\leq T}$ is defined by
$\mathcal{A}_{t,u}^{\underline{\alpha}}:=\int_{0}^{1}w(u,v)\gamma_{t,v}^{\underline{\alpha}}\lambda(\mathrm{d}v),\quad\forall t\in[0,T].$
Thanks to exact law of large numbers (2.7), we can write the aggregate in terms of $\mathbb{E}_{\mathbb{P}}[\gamma_{t,v}^{\underline{\alpha}}$],
$\mathcal{A}_{t,u}^{\underline{\alpha}}$ $=\int_{0}^{1}w(u,v)\mathbb{E}_{\mathbb{P}}[{\gamma_{t,v}^{\underline{\alpha}}}]\lambda(\mathrm{d}v),\quad\forall t\in[0,T].$
We next adopt the following notion of Nash equilibrium.
Definition 6.1.
@@
[…]
@@
The question of the existence of equilibria can be approached using classical arguments from stochastic optimal control, such as the stochastic maximum principle or the dynamic programming principle; see e.g., [82, 19]. We leave this and related issues to future work.